OpenStudy (anonymous):

Jack and Jane are married and both work. However, due to their responsibilities at home, they have decided that they do not want to work over 65 hours per week combined. Jane is paid $12.50 per hour at her job, and Jack is paid $10 per hour at his. Neither of them are paid extra for overtime, but they are allowed to determine the number of hours per week that they wish to work. If they need to make a minimum of $750 per week before taxes, what is the maximum amount of hours that Jack can work per week according to these limits?

OpenStudy (anonymous):

A. 40 B.25 C.30 D.20

OpenStudy (anonymous):

.C

OpenStudy (anonymous):

how? @jayla0428

OpenStudy (anonymous):

wait i think that's wrong

OpenStudy (anonymous):

one sec

OpenStudy (anonymous):

Ok it is .D because since he makes less it wouldn't make sense for him to work more hours then his wife so you do 20 x10=200 and then 45x12.50=562.5 add them together and you get $762.5 which is more then their minimum

OpenStudy (anonymous):

thanks @jayla0428

OpenStudy (anonymous):

you welcome

OpenStudy (jdoe0001):

\(\bf j=\textit{jane's hours}\qquad k=\textit{jack's hours} \\ \quad \\ \textit{total hours they want to work is }65\implies {\color{blue}{ j+k=65}} \\ \quad \\ \textit{she makes }\$12.5\textit{ per hour and he makes }\$10\textit{ per hour}\\ \textit{they need to make at least}\$750\textit{ combined from all hours} \\ \quad \\ \implies {\color{blue}{ 12.5j+10k>=750}} \\ \quad \\ \quad \\ j+k=65\implies j={\color{red}{ 65-k}}\qquad thus \qquad 12.5{\color{red}{ j}}+10k>=750 \\ \quad \\ \implies 12.5{\color{red}{ (65-k)}}+10k>=750\) solve for "k", or Jack's hours